OpenStudy (anonymous):

Solve x + x + 2 + 2x = -2. -2 -1 0 1

OpenStudy (uri):

x+x+2x=-2+2 4x=0 x=0/4 x=0

OpenStudy (lncognlto):

4x = -4 x=-1, not 0